OpenStudy (anonymous):

Have a question about the math coding. I want to write f(x)= 2/x so that the / sign become horizontal. Tried this code f(x)= {2} over {x} but this comes out.

OpenStudy (anonymous):

\[f(x)= {2} \over {x} \]

OpenStudy (lgbasallote):

try \[\text{\[f(x) = \frac{2}{x}\]}\]

OpenStudy (lgbasallote):

tell me if it works or not

OpenStudy (anonymous):

..for what I see on your post, your code doesn't work.

OpenStudy (anonymous):

\[f(x)=\frac{2}{x}\] nice one. better keep this.

OpenStudy (lgbasallote):

lol my code works i intentionally made it not work...here's how it'll look like \[f(x) = \frac{2}{x}\]

OpenStudy (lgbasallote):

i used \text {} so the latex wont function (in case you're interested to learn it)

OpenStudy (anonymous):

how to insert tab within the equation mode?

OpenStudy (lgbasallote):

tab?

OpenStudy (anonymous):

yeah so there is some space.

OpenStudy (anonymous):

I want to write \[a_k =0\] as k -> infinity. in one line.

OpenStudy (lgbasallote):

\[\large a_{k\rightarrow \infty} = 0?\]

OpenStudy (anonymous):

bla bla bla bla (x,y)\in\Sigma\times S

OpenStudy (anonymous):

See it doesn't work this guy can mix text and equation in one line. http://openstudy.com/study#/updates/4ff64f7ee4b01c7be8c957bc

OpenStudy (kinggeorge):

To get \(\LaTeX\) and normal text use \ ( to begin the equation with, and \ ) to end with (make sure to take the space out between the "\" and "(" and ")" ). For example, To write: blarghy blargh \((x,y)\in\Sigma\times S\) blarghidy blurgh. Type: \[\\\(\text{(x,y)\in\Sigma\times S}\)\]

OpenStudy (kinggeorge):

To add spaces in there, use \; for small spaces, \, for extra small spaces, \quad for a larger space, and \qquad for a very large space. Also, if you want to see what others typed to get an equation, right click on the equation select "Show Math As" and click on "TeX commands" There is also a group for this located at http://openstudy.com/study#/groups/LaTeX%20Practicing!%20%3A)